A car traveling 30 m/s will travel m when it uniformly accelerates to 10 m/s with an acceleration of -10 m/s2.
1
Expert's answer
2017-10-25T15:07:07-0400
Kinematic equation
v_f^2-v_i^2=2ad
d=(v_f^2-v_i^2)/2a=((10)^2-(30)^2)/2(-10) =4000 m.
Thus, X=4000.